# Discrete Approximate Circle Bundles

### v2

@@ Abstract
In this paper, we introduce discrete approximate circle bundles, a class of objects designed to serve as the data science analog of circle bundles from algebraic topology. We show that, under appropriate conditions, one can meaningfully and stably identify a discrete approximate circle bundle with an isomorphism class of true circle bundles. We also describe two cohomology invariants which uniquely determine the isomorphism class of a circle bundle, and provide algorithms to compute them given a discrete approximate representative. Finally, we propose a novel methodology for coordinatization and dimensionality reduction of circle bundle data. To illustrate the practical utility and viability of our algorithms, we present applications to both real and synthetic datasets from computer vision (e.g., modeling optical flow).
@@ 1.1 Motivation And Context
In recent years, it has been observed that high-dimensional datasets arising in fields such as computer vision, computational chemistry, and motion tracking often lie near highly non-linear, low-dimensional manifolds [TSL00, RS00, BN03]. These manifolds can exhibit complex geometric and topological structure, motivating the development of a range of techniques for identifying and leveraging such structures in tasks including dimensionality reduction, statistical analysis, pattern recognition, and data visualization [CL06, Sap01]. By modeling the intrinsic manifold geometry, these methods have been shown to improve compression, denoising, and feature extraction in domains where traditional linear techniques fail to capture the underlying structure of the data.
In this paper, we focus on a particular class of datasets which can be modeled as circle bundles. Roughly speaking, a circle bundle is a continuously-varying family of (topological) circles parameterized by another space, called the base space of the bundle. The individual circles are called fibers of the bundle, and the union of the fibers is the total space. The parametrization is given by a projection map $\pi:E\to B$ from the total space to the base space – $E$ is locally homeomorphic to the product $B\times\mathbb{S}^{1}$, but globally the structure may be twisted. For instance, the torus and the Klein bottle can each be characterized as the total space of a circle bundle with base space $\mathbb{S}^{1}$, but the two are not homeomorphic (see Figure 1). Similarly, the unit tangent bundle on the 2-sphere is a circle bundle with base $\mathbb{S}^{2}$, but one can show that there is no bundle isomorphism between this bundle and the product space $\mathbb{S}^{2}\times\mathbb{S}^{1}$.
@@
On the machine learning side, various datasets related to computer vision and graphics have been shown to have circle bundle structures. For instance, in [Ada+20], Adams et al. use local measurements and a feature map to support a torus model for high-contrast optical flow data. Similarly, in [Car+08], the authors propose that a certain space of high-contrast natural image patches is concentrated around a Klein bottle embedded in high dimension. Beyond these examples, datasets arising from 3D objects with rotational symmetries — such as projection images of macromolecules in cryo-electron microscopy — naturally give rise to circle bundle structures, where each fiber corresponds to rotations of the object about a fixed axis [SS14].
The possible circle bundle structures over a metric space $B$ are completely described by a pair of discrete invariants called characteristic classes. In particular, a non-vanishing class indicates the bundle is non-trivial, so there is no hope of constructing a coordinatization for the whole space as a product. Additionally, characteristic classes are invariant under continuous perturbations, so one might hope to be able to stably recover them from a noisy finite sampling. Another key feature of characteristic classes is that they can be computed from local measurements alone, which makes them an ideal tool for working with complex, high-dimensional data not amenable to more direct techniques. Lastly, the characteristic classes of a bundle can be used to construct a so-called bundle map from the total space to a more familiar space which respects topological structure. This observation has recently been leveraged in works such as [Lee+25] and [SP23] to develop non-linear dimensionality reduction techniques based on vector bundles. In this paper, we extend this approach to circle bundles.
@@ 1.2 Our Contributions
    We define notions of discrete approximate circle bundles and local trivializations and show how these relate to other well-studied objects such as discrete approximate Čech cocycles. Building on results of [Tin22] and [SP23], we give conditions under which a discrete approximate circle bundle can be meaningfully identified with a true class of circle bundles.
@@
    We show that every true circle bundle is uniquely characterized by an orientation class and a twisted Euler class, and provide an algorithm for computing these classes from a family of discrete approximate local trivializations. We further introduce a metric on each space of discrete approximate local trivializations and show that two trivializations which are sufficiently close produce the same characteristic classes. In particular, we prove that if a discrete approximate local trivialization is sufficiently close to a true one, the characteristic classes obtained from our algorithm are the same.
@@
    Given a discrete approximate circle bundle over a base space $B$ and an open cover $\mathcal{U}$ for $B$, we introduce a weights filtration on the nerve complex $\mathcal{N}(\mathcal{U})$ and a corresponding notion of persistence for characteristic class representatives. We then propose a scheme for coordinatization and dimensionality reduction which produces a map from the dataset to the total space of the associated bundle $V(2,d)\times_{O(2)}\mathbb{S}^{1}$, where $V(2,d)$ denotes the Stiefel manifold of 2-planes in $\mathbb{R}^{d}$. Our pipeline can be tuned to be compatible with any stage of the weights filtration on $\mathcal{N}(\mathcal{U})$, yielding a point cloud with a different global topology at each stage.
@@
    We provide implementations of our algorithms on several real and synthetic datasets related to imaging and optical flow. In Section 7.1, we use these tools to confirm the torus model proposed in [Ada+20]. A more detailed analysis of our findings will appear in an upcoming companion paper.
@@ 1.3 Prior Work
A number of approaches to local-to-global inference for analyzing geometrically and topologically complex datasets have recently emerged. The basic principle in each case is to model small pieces of the dataset and analyze how these parts glue together to determine the global shape. Given a dataset $X$, a common technique is to choose a feature map $\pi:X\to S$ to some metric space $S$ (e.g., $\mathbb{R}$) and study how the level sets vary. The Mapper pipeline takes this approach to represent a point cloud as a graph with nodes representing local clusters. Similarly, in [SP22] the authors describe a pipeline called FibeRed which uses local PCA together with a feature map $\pi:X\to S$ to obtain a low-dimensional representation of $X$ which is faithful to both local geometry and global topology.
Another aspect of local-to-global analysis is the synchronization of local coordinate systems. Cohomological obstructions to synchronization are explored in works such as [GBM19] using the language of principal $G$-bundles. We note, however, that the theory provided there makes the highly restrictive assumption that the underlying bundle is flat. In [SP23], the authors introduce a theory of discrete approximate Euclidean vector bundles which handles non-flat cases in the context of $O(d)$-synchronization. Building on the language and technology developed in [Tin22], they show that under good conditions, one can uniquely and stably identify discrete approximate Euclidean vector bundles with true ones. They also provide stable algorithms for computing characteristic classes of vector bundles from discrete approximate cocycle representatives.
@@ 1.4 Structure Of Paper
Section 3: Mathematical Background. We review the essential mathematical foundations required for our results, including the general theory of fiber bundles, the classification of circle bundles and cohomology with local coefficients. This section establishes notation and key concepts that will be used throughout the paper.
Section 4: Discrete Approximate Circle Bundles. We introduce the notion of discrete approximate circle bundles and explore their relationship with discrete approximate Čech cocycles and classifying maps as defined in [SP23]. We also discuss how these discrete constructions relate to classical circle bundles.
Section 5: Algorithms for Characteristic Classes. We present algorithms to compute characteristic classes from discrete approximate local trivializations, and prove that our algorithms are stable.
Section 6: Weights Filtration and Persistence. Given a discrete approximate circle bundle over a metric space $B$ and an open cover $\mathcal{U}$ of $B$, we introduce a simplexwise filtration on $\mathcal{N}(\mathcal{U})$ and describe a type of persistent cohomology for characteristic class representatives with respect to this filtration.
Section 7: Coordinatization Pipeline. We provide a pipeline for coordinatization and dimensionality reduction which integrates the Principal Stiefel Coordinates algorithm [Lee+25]. The result is a map from the dataset into a Stiefel manifold which is, in some sense which we explain, compatible with a chosen stage of the associated weights filtration.
Section 8: Implementations and Experiments. We demonstrate our methods on both synthetic and real datasets, applying the sparse circular coordinates algorithm [Per21] to obtain discrete approximate local trivializations and showcasing the resulting computations of characteristic classes and coordinatization.
@@ 2 Preliminaries
2.1 Theory Of Fiber Bundles
In this section, we briefly review the key aspects of the theory of fiber bundles which are relevant for us. Our exposition emphasizes a computational perspective. For additional details, see [Hus94].
@@ 2.1 Theory Of Fiber Bundles
Any such map $\varphi$ is called a local trivialization of $\xi$. An open cover $\mathcal{U}$ of $B$ is called a trivializing cover for $\xi$ if there exists a local trivialization of $\xi$ over each $U_{j}\in\mathcal{U}$. Two bundles $\xi=(E,\pi,B)$ and $\xi^{\prime}=(E^{\prime},\pi^{\prime},B)$ are isomorphic if there exists a homeomorphism $f:E\to E^{\prime}$ such that $\pi^{\prime}\circ f=\pi$. Let $\text{Bun}_{F}(B)$ denote the set of isomorphism classes.
Transition Maps: Given a fiber bundle $\xi=(E,\pi,B)$ with fiber $F$, an open cover $\mathcal{U}$ of $B$ and a family $\{\varphi_{j}:\pi^{-1}(U_{j})\to U_{j}\times F\}_{j\in J}$ of local trivializations subordinate to $\mathcal{U}$, there exist unique continuous maps $\Omega_{jk}:U_{j}\cap U_{k}\to\text{Homeo}(F)$ such that $\varphi_{j}\circ\varphi_{k}^{-1}(b,v)=(b,\Omega_{jk}(b)(v))$ for any $(b,v)\in(U_{j}\cap U_{k})\times F$. The maps $\{\Omega_{jk}\}_{(jk)\in\mathcal{N}(\mathcal{U})}$ are called the transition maps associated with the family of local trivializations. Note that the transition functions associated with any family of local trivializations must satisfy the cocycle condition:
$\Omega_{jk}(b)\circ\Omega_{kl}(b)=\Omega_{jl}(b)\hskip 28.45274pt(b\in U_{jkl}\;,\;(jkl)\in\mathcal{N}(\mathcal{U}))$ (1)
@@
Part A: Let $B$ and $F$ be topological spaces, and let $G$ be a topological group with a continuous left action on $F$. Given an open cover $\mathcal{U}=\{U_{j}\}_{j\in J}$ of $B$ and a collection $\{\Omega_{jk}:U_{j}\cap U_{k}\to G\}_{(jk)\in\mathcal{N}(\mathcal{U})}$ satisfying the cocycle condition, there exists a fiber bundle $\pi:E\to B$ with fiber $F$ and structure group $G$ that is trivializable over $\mathcal{U}$ with transition maps $\{\Omega_{jk}\}_{(jk)\in\mathcal{N}(\mathcal{U})}$.
Part B: Suppose $\xi^{\prime}=(E^{\prime},\pi^{\prime},B^{\prime})$ is another fiber bundle with fiber $F$ and structure group $G$ which is trivializable over $\mathcal{U}$ with transition maps $\{\Omega_{jk}^{\prime}\}_{(jk)\in\mathcal{N}(\mathcal{U})}$. If the action of $G$ on $F$ is faithful, then $\xi\cong\xi^{\prime}$ if and only if there exist maps $\{\mu_{j}:U_{j}\to G\}_{j\in J}$ such that $\Omega_{jk}^{\prime}=\mu_{j}^{-1}\Omega_{jk}\mu_{k}$ for all $(jk)\in\mathcal{N}(\mathcal{U})$.
@@
Associated Bundles: Let $\xi=(E,\pi,B)$ be a fiber bundle with fiber $F$ and structure group $G$, and let $F^{\prime}$ be a left $G$-space. The associated bundle to $\xi$ with fiber $F^{\prime}$ is the bundle $\xi^{\prime}=(E\times_{G}F^{\prime},\pi^{\prime},B)$ whose existence is guaranteed by the fiber bundle construction lemma (using the transition maps from any trivialization of $\xi$). The special case of $F^{\prime}=G$ with left action by group multiplication is called the principal $G-\textbf{bundle}$ associated with $\xi$. Note that for any space $B$, the associated bundle construction gives a bijective correspondence between fiber bundles with fiber $F$ and principal $\text{Homeo}(F)$-bundles over $B$.

### v3

# Brad Turow, Jose A. Perea

@@ Abstract
In this paper, we introduce discrete approximate circle bundles, a class of objects designed to serve as the data science analog of circle bundles from algebraic topology. We show that, under appropriate conditions, one can meaningfully and stably identify a discrete approximate circle bundle with an isomorphism class of true circle bundles. We also describe two cohomology invariants which uniquely determine the isomorphism class of a circle bundle, and provide algorithms to compute them given a discrete approximate representative. Finally, we propose a novel methodology for coordinatization and dimensionality reduction of circle bundle data. To illustrate the practical utility and viability of our algorithms, we present applications to both real and synthetic datasets from computer vision (e.g., modeling optical flow). The paper is accompanied by an open-source software package, with full documentation and tutorials, enabling reproducible implementation of the proposed algorithms and experiments, including those used to generate the figures in this paper
@@ 1.1 Motivation And Context
It has been observed in recent years that high-dimensional datasets arising in fields such as computer vision, computational chemistry, and motion tracking often lie near highly non-linear, low-dimensional manifolds [29, 21, 2]. These manifolds can exhibit complex geometric and topological structure, motivating the development of a range of techniques for identifying and leveraging such structures in tasks including dimensionality reduction, statistical analysis, pattern recognition, and data visualization [6, 23]. By modeling the intrinsic manifold geometry, these methods have been shown to improve compression, denoising, and feature extraction in domains where traditional linear techniques fail to capture the underlying structure of the data [15].
In practice, complex non-linear structure in large, high-dimensional datasets is often difficult to observe directly; it is therefore essential to have techniques for inferring global structure from purely local computations on small portions of the data. In this paper, we focus on a particular class of datasets which can be modeled as circle bundles. Roughly speaking, a circle bundle is a continuously-varying family of (topological) circles parameterized by another space, called the base space of the bundle. The individual circles are called the fibers of the bundle, and the union of the fibers is the total space. The parametrization is given by a projection map $\pi:E\to B$ from the total space to the base space; in particular $E$ is locally homeomorphic to a product $U\times\mathbb{S}^{1}$, for $U\subseteq B$ open, but globally the structure may be twisted. For instance, the torus and the Klein bottle can each be characterized as the total space of a circle bundle with base space $\mathbb{S}^{1}$, but the two are not homeomorphic (see Figure 1). Similarly, the 3-manifold $SO(3)$ admits a non-trivial circle bundle structure over $\mathbb{S}^{2}$. Indeed, consider the map $\pi:SO(3)\to\mathbb{S}^{2}$ which assigns an (oriented) axis of rotation to each element of $SO(3)$; the fibers of this map are copies of $SO(2)\cong\mathbb{S}^{1}$, but one can show that $SO(3)\ncong\mathbb{S}^{2}\times\mathbb{S}^{1}$.
@@
On the machine learning side, various datasets related to computer vision and graphics have been shown to have circle bundle structures. For instance, in [1], Adams et al. use local measurements and a feature map to support a torus model for high-contrast optical flow data. Similarly, in [5], it is shown that frequently-occurring high-contrast patches from natural images concentrate around a Klein bottle embedded in a high-dimensional Euclidean space. It is worth noting that in each of these contexts, a direct persistent homology computation fails to satisfactorily confirm the topology of the proposed underlying model. Figure 2 shows an illustration of how this failure can occur, and shows how the local-to-global methodologies we propose in this paper successfully recover the underlying structure.
Refer to caption (a) Persistence diagrams of the original toroidal dataset $X$ Refer to caption (b) $X$ colored by the circular coordinate derived from its most persistent 1-dimensional cohomology class [20] Refer to caption (c) Fiberwise angle resulting from trivializing the circle bundle over $\mathbb{S}^{1}$ given by (b) Refer to caption
Figure 2: A noisy sample $X$ from a torus whose fiber radius continuously oscillates. The persistence diagrams show only a single significant class in dimension 1 (recall that the non-zero Betti numbers for the torus are $\beta_{0}=1$, $\beta_{1}=2$, $\beta_{2}=1$). Figure (b) shows the dataset colored according to circular coordinates computed with DREiMac [19]. Figure (c) shows the same dataset colored according to the fiber coordinates computed using the software package accompanying this paper.
Beyond these examples, datasets arising from 3D objects with rotational symmetries — such as projection images of macromolecules in cryo-electron microscopy — naturally give rise to circle bundle structures, where each fiber corresponds to rotations of the object about a fixed axis [26]. Here the underlying model is typically a 3-manifold such as $SO(3)$, and a direct persistence computation to capture the topology is simply intractable. On the other hand, one can infer the underlying manifold from purely local computations by leveraging the bundle structure, as we demonstrate with a synthetic dataset of 3D density functions in Section 7.3.
The different possible circle bundle structures over a topological space $B$ are completely described, up to isomorphism, by a pair of discrete invariants called characteristic classes. In particular, a non-vanishing class indicates the bundle is non-trivial, so there is no hope of constructing a coordinatization for the whole space as a product. Additionally, characteristic classes are invariant under continuous perturbations, so one might hope to be able to stably recover them from a noisy finite sampling. Another key feature of characteristic classes is that they can be computed from local measurements alone, which makes them suitable for working with complex, high-dimensional data not amenable to more direct techniques. Lastly, the characteristic classes of a bundle can be used to construct a so-called bundle map from the total space to a more familiar space which respects topological structure. This observation has recently been leveraged in works such as [14] and [24] to develop non-linear dimensionality reduction techniques based on vector bundles. In this paper, we extend this approach to circle bundles.
@@ 1.2 Our Contributions
    We define notions of discrete approximate circle bundles (Definition 3.8) in terms of approximate local trivializations (Definition 3.6), and show how these relate to other well-studied objects such as discrete approximate Čech cocycles (Proposition 3.15). Building on results of [30] and [24], we give conditions under which a discrete approximate circle bundle can be meaningfully identified with a true class of circle bundles (Theorem 3.42).
@@
    We show that every true circle bundle is uniquely characterized up to isomorphism by an orientation class and a twisted Euler class (Theorem 2.12), and provide an algorithm for computing these classes from a system of discrete approximate local trivializations (Algorithm 1). We further introduce a metric on each space of discrete approximate local trivializations and show that two trivializations which are sufficiently close produce the same characteristic classes (Corollary 4.3). In particular, we prove that if a discrete approximate local trivialization is sufficiently close to a true one, the characteristic classes obtained from our algorithm are the same (Corollary 4.5).
@@
    Given a discrete approximate circle bundle over a base space $B$ and an open cover $\mathcal{U}$ for $B$, we introduce a weights filtration on the nerve complex $\mathcal{N}(\mathcal{U})$ and a corresponding notion of persistence for characteristic class representatives (Section 5.1). We then propose a scheme for coordinatization and dimensionality reduction which produces a map from the dataset to the total space of the associated bundle $V(2,d)\times_{O(2)}\mathbb{S}^{1}$, where $V(2,d)$ denotes the Stiefel manifold of 2-planes in $\mathbb{R}^{d}$ (Section 6). Our pipeline is designed to be compatible with any stage of the weights filtration on $\mathcal{N}(\mathcal{U})$, yielding multi-scale representations and coordinatizations of the data consistent with its multi-scale global topology.
@@
    We demonstrate implementations of our algorithms on several real and synthetic datasets related to imaging and optical flow. In Section 7.1, we use these tools to confirm the torus model proposed in [1]. A more detailed analysis of our findings can be found in our companion paper [32].
  5. 5.
    We provide an open-source software package implementing the algorithms introduced in this paper, available at https://github.com/bradturow/Circle_Bundles, with full documentation and tutorials hosted at https://circle-bundles.readthedocs.io
@@ 1.3 Prior Work
Recently, there has been an emergence of new approaches in local-to-global inference for analyzing geometrically and topologically complex datasets. The basic principle in each case is to model small pieces of the dataset and analyze how these parts glue together to determine the global shape. Given a dataset $X$, a common technique is to choose a feature map $\pi:X\to S$ to some metric space $S$ (e.g., $\mathbb{R}$) and study how the level sets vary. The Mapper pipeline takes this approach to represent a point cloud as a graph with nodes representing local clusters [27]. Similarly, in [25] the authors describe a pipeline called FibeRed which uses linear dimensionality reduction — e.g., via PCA or ISOMAP — on the fibers of a feature map $\pi:X\to S$ to a Riemannian manifold $S$, to obtain a low-dimensional representation of $X$ which captures both local geometry and global topology.
Another aspect of local-to-global analysis is the synchronization of local coordinate systems. Cohomological obstructions to synchronization are explored in works such as [9] using the language of principal $G$-bundles. We note, however, that the theory provided there makes the restrictive assumption that the underlying bundle is flat (which is not the case in most of the applications we consider). In [24], the authors introduce a theory of discrete approximate Euclidean vector bundles which handles non-flat cases in the context of $O(d)$-synchronization. Building on the language and techniques developed in [30], they show that under appropriate conditions, one can uniquely and stably identify discrete approximate Euclidean vector bundles with true ones. They also provide stable algorithms for computing characteristic classes of vector bundles from discrete approximate cocycle representatives.
@@ 1.4 Structure Of Paper
Section 2: Mathematical Preliminaries. We review the essential mathematical elements required for our results, including the general theory of fiber bundles, the classification of circle bundles and cohomology with local coefficients. This section establishes notation and key concepts that will be used throughout the paper.
Section 3: Discrete Approximate Circle Bundles. We introduce our proposed notion of discrete approximate circle bundles and explore its relationship with discrete approximate Čech cocycles and classifying maps as defined in [24]. We also discuss how these discrete constructions relate to classical circle bundles.
Section 4: Algorithms for Characteristic Classes. We present algorithms to compute characteristic classes from discrete approximate local trivializations, and prove that our algorithms are robust to input perturbations (i.e., stable).
Section 5: Weights Filtration and Persistence. Given a discrete approximate circle bundle over a metric space $B$ and an open cover $\mathcal{U}$ of $B$, we introduce a filtration on the nerve complex $\mathcal{N}(\mathcal{U})$ and describe a type of persistent cohomology for characteristic class representatives with respect to this filtration.
Section 6: Coordinatization Pipeline. We provide a pipeline for coordinatization and dimensionality reduction of circle bundle data, which integrates the Principal Stiefel Coordinates algorithm [14]. The result is a function from the dataset into a Stiefel manifold which is, in some sense which we explain, compatible with a chosen stage of the associated weights filtration.
Section 7: Implementations and Experiments. We demonstrate our methods on both synthetic and real datasets, applying the sparse circular coordinates algorithm [20] to obtain discrete approximate local trivializations and showcasing the resulting computations of characteristic classes and coordinatization.
@@ 2 Preliminaries
2.1 Fiber Bundles
In this section, we briefly review the key aspects of the theory of fiber bundles which are relevant for us. Our exposition emphasizes a computational perspective. For additional details, see [12].
@@ 2.1 Fiber Bundles
Any such map $\varphi$ is called a local trivialization of $\xi$. An open cover $\mathcal{U}$ of $B$ is called a trivializing cover for $\xi$ if there exists a local trivialization of $\xi$ over each $U\in\mathcal{U}$. Two bundles $\xi=(E,\pi,B)$ and $\xi^{\prime}=(E^{\prime},\pi^{\prime},B)$ are isomorphic if there exists a homeomorphism $f:E\to E^{\prime}$ such that $\pi^{\prime}\circ f=\pi$. Let $\text{Bun}_{F}(B)$ denote the set of isomorphism classes of fiber bundles over $B$ with fiber $F$.
Definition 2.2.
Let $G$ be a topological group. A principal $G$-bundle is a fiber bundle $\xi=(E,\pi,B)$ with fiber $G$, satisfying the following properties: $E$ has a free right $G$-action so that $\pi(eg)=\pi(e)$ for all $(e,g)\in E\times G$ — i.e., the action is fiber-preserving — and which is transitive when restricted to each fiber $\pi^{-1}(b)$, $b\in B$. Moreover the local trivializations $\varphi:\pi^{-1}(U_{b})\to U_{b}\times G$ can be chosen to be $G$-equivariant, where the right $G$-action on $U_{b}\times G$ is group multiplication in $G$. Two principal $G$-bundles $(E,\pi,B)$ and $(E^{\prime},\pi^{\prime},B)$ are isomorphic (as principal $G$-bundles) if there exists a $G$-equivariant homeomorphism $f:E\to E^{\prime}$ so that $\pi^{\prime}\circ f=\pi$. Let $\text{Prin}_{G}(B)$ denote the set of isomorphism classes of principal $G$-bundles over $B$.
Transition Maps: Given a fiber bundle $\xi=(E,\pi,B)$ with fiber $F$, a trivializing cover $\mathcal{U}=\{U_{j}\}_{j\in J}$ of $B$ and a family $\{\varphi_{j}:\pi^{-1}(U_{j})\to U_{j}\times F\}_{j\in J}$ of local trivializations subordinate to $\mathcal{U}$, there exist unique continuous maps $\Omega_{jk}:U_{j}\cap U_{k}\to\text{Homeo}(F)$ such that $\varphi_{j}\circ\varphi_{k}^{-1}(b,v)=(b,\Omega_{jk}(b)(v))$ for any $(b,v)\in(U_{j}\cap U_{k})\times F$. Here $\text{Homeo}(F)$ denotes the set of homeomorphisms from $F$ to $F$, endowed with the compact-open topology. The maps $\{\Omega_{jk}\}_{(jk)\in\mathcal{N}(\mathcal{U})}$, indexed by the 1-simplices of the nerve of $\mathcal{U}$, are called the transition maps associated with the family of local trivializations. Note that the transition functions associated with any family of local trivializations must satisfy the cocycle condition:
$\Omega_{jk}(b)\circ\Omega_{kl}(b)=\Omega_{jl}(b)\hskip 28.45274pt(b\in U_{j}\cap U_{k}\cap U_{l}\;,\;(jkl)\in\mathcal{N}(\mathcal{U}))$ (1)
@@
Part A: Let $B$ and $F$ be topological spaces, and let $G$ be a topological group with a continuous left action on $F$. Given an open cover $\mathcal{U}=\{U_{j}\}_{j\in J}$ of $B$ and a collection $\{\Omega_{jk}:U_{j}\cap U_{k}\to G\}_{(jk)\in\mathcal{N}(\mathcal{U})}$ of continuous maps satisfying the cocycle condition (1), there exists a fiber bundle $\pi:E\to B$ with fiber $F$ and structure group $G$ that is trivializable over $\mathcal{U}$ with transition maps $\{\Omega_{jk}\}_{(jk)\in\mathcal{N}(\mathcal{U})}$ [12, Chapter 5, Theorem 3.2].
Part B: Suppose $\xi^{\prime}=(E^{\prime},\pi^{\prime},B)$ is another fiber bundle with fiber $F$ and structure group $G$ (endowed with the same left-action) which is trivializable over $\mathcal{U}$ with transition maps $\{\Omega_{jk}^{\prime}\}_{(jk)\in\mathcal{N}(\mathcal{U})}$. If the action of $G$ on $F$ is faithful, then $\xi\cong\xi^{\prime}$ if and only if there exist maps $\{\mu_{j}:U_{j}\to G\}_{j\in J}$ such that $\Omega_{jk}^{\prime}=\mu_{j}^{-1}\Omega_{jk}\mu_{k}$ for all $(jk)\in\mathcal{N}(\mathcal{U})$ [12, Chapter 5, Theorem 2.7].
@@
Associated Bundles: Let $\xi=(E,\pi,B)$ be a fiber bundle with fiber $F$ and structure group $G$. Let $\mathcal{U}=\{U_{j}\}_{j\in J}$ be a trivializing cover for $\xi$ and let $\{\Omega_{jk}:U_{j}\cap U_{k}\to G\}_{(jk)\in\mathcal{N(U)}}$ be the transition maps of a local trivializing of $\xi$ over $\mathcal{U}$. If $F^{\prime}$ is a left $G$-space, then Part A of the Fiber Bundle Construction Lemma defines a fiber bundle $\xi^{\prime}$ over $B$ with fiber $F^{\prime}$. The association $\xi\mapsto\xi^{\prime}$ is well-defined up to isomorphism of fiber bundles, and we call $\xi^{\prime}$ the associated bundle to $\xi$ with fiber $F^{\prime}$. The special case $F^{\prime}=G$ with left action by group multiplication is called the principal $G-\textbf{bundle}$ associated to $\xi$. When $\xi=(E,\pi,B)$ is a principal $G$-bundle and $F^{\prime}$ is a left $G$-space, one can construct $\xi^{\prime}$ directly without appealing to transition functions or isomorphism classes as follows: The multiplication $\cdot:(E\times F^{\prime})\times G\to E\times F^{\prime}$ given by $(e,v^{\prime})\cdot g=(eg\,,\,g^{-1}*v^{\prime})$ defines a right $G$-action on $E\times F^{\prime}$, and the orbit space $E\times_{G}F^{\prime}$ has a map $\pi^{\prime}:E\times_{G}F^{\prime}\to B$ given by $\pi^{\prime}([e,v^{\prime}])=\pi(e)$. One can show that $\xi^{\prime}=(E\times_{G}F^{\prime},\pi^{\prime},B)$ is a fiber bundle with fiber $F^{\prime}$. Finally, we note that for any space $B$, the associated bundle construction gives a bijective correspondence between $\text{Bun}_{F}(B)$ and $\text{Prin}_{\text{Homeo}(F)}(B)$, where $F$ is understood to be endowed with the tautological $\text{Homeo}(F)$-action by evaluation.